# Haotong Yang    Zitong Wang    Shijia Kang    Siqi Yang    Wenkai Yu    Xu Niu    Yike Sun    Yi Hu    Zhouchen Lin    Muhan Zhang

### v1

@@ A.8 Error type breakdown and case study of ProofRM
In this section, we provide some detailed analysis of our behavior ProofRM. First, we find that our model has been successfully learn to recognize some deduction error in the proof, which is hard to recognize for the untrained baseline. Below we show some examples:
@@
We can see from this example above that the checked proof contains fatal deduction errors (highlighted red) in both cases, being ambiguous on the reason of why should $n$ exists satisfying the given criterion. The checked proof relies upon intuitive characterizations of $\epsilon$, which is correctly identified in the green part of the output of our model.
@@
Example of learned erroneous logic flow Question. The Y2K Game is played on a $1\times 2000$ grid as follows. Two players in turn write either an S or an O in an empty square. The first player who produces three consecutive boxes that spell SOS wins. If all boxes are filled without producing SOS then the game is a draw. Prove that the second player has a winning strategy. Answer. The main insight is that a construct of the form $S\;\square\;\square\;S$ (here the $\square$ is blank) will kill any player which plays inside it. We call this a trap accordingly. The actual strategy is as follows: the core strategy of the second player is to win via producing traps. A trap is a structure of the form $S\;\square\;\square\;S$. If any player is forced to place a letter in either of the spaces, then the opponent can immediately complete the structure SOS and thus win this game. Specifically, the second player prioritizes constructing such traps when deciding the current action through the following: Select a grid where an S is already present, and place another S on the third blank grid away from the current grid (i.e. leaving two empty grids in between the two S), while satisfying the following requirement: (1) the two square in between the two S are empty; (2) Placing S in this square will not form an SOS immediately; (3) This action will not lead to the immediate defeat of the player. If no such traps can be constructed, then she will adopt the safe playing method as described in the upcoming claim. As there are a total of 2000 squares in this grid, she can always construct a trap. Once the trap is constructed, the first player will be forced to write a letter in on of the squares (since all squares will be eventually filled up), then the second player only needs to fill in the correct letter in the remaining space (if the opponent writes O in the first empty space then she write S to form SOS; otherwise she write O to form SOS) and that will lead to the victory of the second player. Claim The second player always has a move which prevents her from losing. Proof Since there are an odd number of empty squares at the start of the second player’s turn, there must be a square which is bordered by either two filled or two empty squares. The second player can then play $O$ in this square, which is always safe. Together these two claims finish the problem. Baseline output. The provided proof is intended to demonstrate that the second player has a winning strategy in the Y2K Game on a 1x2000 grid. The proof consists of an initial explanation of a trap strategy and a formal claim with its proof. After evaluating the proof against the criteria for logical soundness, computational accuracy, structural completeness, and notational rigor, it is concluded that the proof is correct. ### Logical Soundness - **Inferences and Justifications:** The proof implies logical flow, with the trap strategy as a key mechanism for P2 to force a win. The trap creation ($S\;\square\;\square\;S$) and safety conditions are well-justified, and the move-by-move deductive reasoning for SOS completions appears sound. The claim about P2 always having a safe move is supported by the existence of a square with both neighbors same (due to E odd and S large), and playing O there prevents immediate loss. - **Lemmas/Theorems:** The proof implicitly uses properties of a linear grid, and the claim relies on the parity of E, which holds since S=2000 even and E odd at starts of P2 turns. Conditions for the trap are satisfied as per the strategy description. - **Gaps/Minor Issues:** There is ambiguity in the definition of ”bordered by two squares” for end squares, but since S is large and even with E odd, the existence of such a square (with both neighbors same) is a standard result for large grids and does not affect the core argument. The trap strategy is not fully detailed, but the explanation is sufficient for understanding. ### Computational Accuracy - The grid size S=2000 is used, but no numerical calculations are involved in the proof, so computational accuracy is not applicable. The validity of SOS completions and neighbors checks (e.g., $S\;\square\;\square\;S$) are contextually correct based on game mechanics. ### Structural Completeness - The proof fully addresses the problem: P2 has a winning strategy either by creating a trap or using a safe move as per the claim. - The flow is organized, starting with the trap concept, then the formal claim, with the proof for the claim. No essential elements of the game or strategy are omitted. ### Notational Rigor - Mathematical notation is used (e.g., $S\;\square\;\square\;S$, E for empty), but it is explained clearly, with no ambiguous or non-standard symbols. Definitions (trap, bordered) are implicit, but standard English is used in the proof steps. Notational rigor is acceptable given the expository style. ### Conclusion The proof rigorously supports the claim that the second player has a winning strategy by combining the trap mechanism and the safe move guarantee from the claim. The existence condition for the safe move holds due to the grid size and E odd, and the reasoning is logically sound. ### True ProofRM output. The proof attempts to establish that the second player can force a win by constructing traps (specifically, sequences $S\;\square\;\square\;S$) and by ensuring always having a safe move, such as playing O in certain positions. **(1) Logical Soundness** - **Inferences and Definitions:** The proof introduces undefined terms like ”trap” and a bordering condition (squares bordered by two filled or two empty squares). These are not rigorously defined or justified from standard game theory. The inference that a trap ($S\;\square\;\square\;S$) can force a win is not validated; it is stated as a strategy based on the second player’s ability to place S three away from existing S, but this depends on prior positions and may not always be applicable or avoid immediate loss. For example, placing S three squares away might not be possible if the grid is fragmented or if other conditions are violated (e.g., existing S with closer neighbors). - **Safe Move Claim:** The claim that a square is bordered by two filled or two empty squares due to an odd number of empty squares lacks basis. The number of empty squares being odd at the second player’s turn is unrelated to the bordering condition. The absence of such a border might not prevent a safe move, and the safe move is defined as O without justification (typically in the second player’s turn, playing O might not always be safe due to potential SOS threats from adjacent cells). - **Argument Structure:** The proof combines two potential strategies (traps and safe moves), but their integrations are vague. The safe move claim is not proven logically; it relies on an unverified implicit assumption, similar to the trapping strategy. This creates gaps in reasoning, and special-case analysis for safe moves cannot replace rigorous proof. - **Logical Fallacies:** The proof claims that traps ”kill” the opponent, but this is not demonstrated; it is asserted without showing that all first player responses are covered. The safe move relies on misstated bordering; for instance, corner cells have fewer neighbors, but the grid is linear, making the border condition ill-defined for edges. **(2) Computational Accuracy** - **Mathematical Calculations:** There are no significant calculations, so computational accuracy is not an issue, but this implies mathematical insight is missing. **(3) Structural Completeness** - **Full Coverage:** The proof does not address cases where the first player might prevent trap construction or exploit weaknesses in the safe move strategy early on. It assumes the second player can always build traps or perform the safe move, but no upper bound is established. The problem requires a winning strategy, but no complete game tree or position is analyzed. - **Proof Flow:** The proof jumps between non-verified claims and undefined terms. The safe move is presented as evidence for the main claim, but it is only part of the argument; the entire strategy is incomplete. **(4) Notational Rigor** - **Mathematical Notation:** The proof uses non-standard language like ”trap” and ambiguous descriptions such as ”bordered by two empty squares” applied to squares with an unclear number of neighbors. Terms like $\square$” and variable ”S” are not formally defined. - **Justification and Definitions:** Key concepts like SOS completion are not fully substantiated. The claim of a safe O move lacks proof of time-independent security (e.g., it might lead to SOS if the grid has other S around). The proof is logically flawed and incomplete, with gaps in reasoning and undefined elements. ### False
@@
We can observe from this example above that the original proof lacks necessary clarification to make the winning strategy viable and logically sound, relying on instinctive arguments when providing evidence to why the strategy leads to a winning situation. Our model is able to correctly discern this erroneous logic flow, as seen in the green highlighted part.
@@
Trying to provide a statistical analysis of which type of error can be better recognized by our ProofRM, we sample 97 QA pairs and manually categorized them into three types: misusing unjustified assumptions, incomplete logic flow, and making up non-existent conclusions.
@@
The category comes from our idea that a logic step can be roughly modeled as a syllogism. So we categorize the errors based on the three key elements of a syllogism: premises, logic flow, and conclusion. The first error type corresponds to issues with premises, the second to issues with logic flow, and the third to issues with conclusions. We find most errors are due to incomplete logic flow, which indicates that the model fails to chain the reasoning steps rigorously. The Table 9 shows the results.
@@
Let’s analyze the results in the table. Notice that the performance of base model is even lower than random guess (50%), suggesting the model easily believes incorrect proofs. ProofRM improves accuracy across these error types, with higher gains in the second and third categories, where the base models shows weaker performance. It shows that our RL stage indeed helps our model to avoid the easy-believing behavior and find these errors.
@@ A.9 Ablation for balanced token weighting
In this section, we demonstrate the ablation results for balanced token weighting, giving a quantitative analysis on the resulting stability advantages caused by the introduction of this trick. We display the output length and overall reward of different token weighting strategies, namely GRPO-style, DAPO-style and balanced token weighting, throughout the training procedure in Figure 5 and Figure 6 respectively. The figures demonstrate the effectiveness of balanced-token-weighting. Although the reward for GRPO-style token weighting is increasing with training steps, upon closer inspection of the output we can see that the outputs contain overly short or even no reasoning, indicating the existence of behavior degeneration.
Figure 5: Output length comparison of different token weighting strategies
Figure 6: Reward comparison of different token weighting strategies
@@ A.10 Experiment results on other datasets
In this section we demonstrate the performance of our ProofRM on other evaluation sets, including both those readily provided by the community and those that we constructed ourselves. Table 10 displays the performance of ProofRM with differing model sizes on ProofBench (Ma et al., 2025). We can observe that our ProofRM outperform baseline models of all sizes.
@@
Table 11: Results of ProofRM on mathematical proofs of different styles.
@@ A.11 Ablation on data diversity dimensions
Table 12: Ablation results on different data diversity dimensions. w/o Multi-sourced Dataset indicates that only samples from OlympiadBench is kept. w/o Extended LLM Usage indicates that only samples from Deepseek-R1 is kept, w/o Diverse Prompt Method indicates that only samples using rephrase is kept. Here the accuracy is evaluated on the our test-set.
@@
We may observe from Table 12 that each dimension contributes to the diverse data quantity and also the final performance. Specifically, the diversity introduced by the ”prompt method” is the key to train a generalized model, where the model without this diversity shows worse accuracy than its base (69.9% v.s. 71.6%). The results also indicates the necessity and advantages of scaling up the dataset, further justifying our pipeline to produce data in a scalable way.

### v2

@@ A.8 Error type breakdown and case study of ProofRM
In this section, we provide some detailed analysis of our behavior ProofRM. First, we find that our model has successfully learned to recognize some deduction error in the proof, which is hard to recognize for the untrained baseline. Below we show some examples:
@@
We can see from this example above that the checked proof contains fatal deduction errors (highlighted red) in both cases, being ambiguous on the reason why $n$ should exists satisfying the given criterion. The checked proof relies upon intuitive characterizations of $\epsilon$, which is correctly identified in the green part of the output of our model.
@@
Example of learned erroneous logic flow Question. The Y2K Game is played on a $1\times 2000$ grid as follows. Two players in turn write either an S or an O in an empty square. The first player who produces three consecutive boxes that spell SOS wins. If all boxes are filled without producing SOS then the game is a draw. Prove that the second player has a winning strategy. Answer. The main insight is that a construct of the form $S\;\square\;\square\;S$ (here the $\square$ is blank) will kill any player which plays inside it. We call this a trap accordingly. The actual strategy is as follows: the core strategy of the second player is to win via producing traps. A trap is a structure of the form $S\;\square\;\square\;S$. If any player is forced to place a letter in either of the spaces, then the opponent can immediately complete the structure SOS and thus win this game. Specifically, the second player prioritizes constructing such traps when deciding the current action through the following: Select a grid where an S is already present, and place another S on the third blank grid away from the current grid (i.e. leaving two empty grids in between the two S), while satisfying the following requirement: (1) the two square in between the two S are empty; (2) Placing S in this square will not form an SOS immediately; (3) This action will not lead to the immediate defeat of the player. If no such traps can be constructed, then she will adopt the safe playing method as described in the upcoming claim. As there are a total of 2000 squares in this grid, she can always construct a trap. Once the trap is constructed, the first player will be forced to write a letter in on of the squares (since all squares will be eventually filled up), then the second player only needs to fill in the correct letter in the remaining space (if the opponent writes O in the first empty space then she write S to form SOS; otherwise she write O to form SOS) and that will lead to the victory of the second player. Claim The second player always has a move which prevents her from losing. Proof Since there are an odd number of empty squares at the start of the second player’s turn, there must be a square which is bordered by either two filled or two empty squares. The second player can then play $O$ in this square, which is always safe. Together these two claims finish the problem. Baseline output. The provided proof is intended to demonstrate that the second player has a winning strategy in the Y2K Game on a 1x2000 grid. The proof consists of an initial explanation of a trap strategy and a formal claim with its proof. After evaluating the proof against the criteria for logical soundness, computational accuracy, structural completeness, and notational rigor, it is concluded that the proof is correct. ### Logical Soundness - **Inferences and Justifications:** The proof implies logical flow, with the trap strategy as a key mechanism for P2 to force a win. The trap creation ($S\;\square\;\square\;S$) and safety conditions are well-justified, and the move-by-move deductive reasoning for SOS completions appears sound. The claim about P2 always having a safe move is supported by the existence of a square with both neighbors same (due to E odd and S large), and playing O there prevents immediate loss. - **Lemmas/Theorems:** The proof implicitly uses properties of a linear grid, and the claim relies on the parity of E, which holds since S=2000 even and E odd at starts of P2 turns. Conditions for the trap are satisfied as per the strategy description. - **Gaps/Minor Issues:** There is ambiguity in the definition of ”bordered by two squares” for end squares, but since S is large and even with E odd, the existence of such a square (with both neighbors same) is a standard result for large grids and does not affect the core argument. The trap strategy is not fully detailed, but the explanation is sufficient for understanding. ### Computational Accuracy - The grid size S=2000 is used, but no numerical calculations are involved in the proof, so computational accuracy is not applicable. The validity of SOS completions and neighbors checks (e.g., $S\;\square\;\square\;S$) are contextually correct based on game mechanics. ### Structural Completeness - The proof fully addresses the problem: P2 has a winning strategy either by creating a trap or using a safe move as per the claim. - The flow is organized, starting with the trap concept, then the formal claim, with the proof for the claim. No essential elements of the game or strategy are omitted. ### Notational Rigor - Mathematical notation is used (e.g., $S\;\square\;\square\;S$, E for empty), but it is explained clearly, with no ambiguous or non-standard symbols. Definitions (trap, bordered) are implicit, but standard English is used in the proof steps. Notational rigor is acceptable given the expository style. ### Conclusion The proof rigorously supports the claim that the second player has a winning strategy by combining the trap mechanism and the safe move guarantee from the claim. The existence condition for the safe move holds due to the grid size and E odd, and the reasoning is logically sound. ### True ProofRM output. The proof attempts to establish that the second player can force a win by constructing traps (specifically, sequences $S\;\square\;\square\;S$) and by ensuring always having a safe move, such as playing O in certain positions. **(1) Logical Soundness** - **Inferences and Definitions:** The proof introduces undefined terms like ”trap” and a bordering condition (squares bordered by two filled or two empty squares). These are not rigorously defined or justified from standard game theory. The inference that a trap ($S\;\square\;\square\;S$) can force a win is not validated; it is stated as a strategy based on the second player’s ability to place S three away from existing S, but this depends on prior positions and may not always be applicable or avoid immediate loss. For example, placing S three squares away might not be possible if the grid is fragmented or if other conditions are violated (e.g., existing S with closer neighbors). - **Safe Move Claim:** The claim that a square is bordered by two filled or two empty squares due to an odd number of empty squares lacks basis. The number of empty squares being odd at the second player’s turn is unrelated to the bordering condition. The absence of such a border might not prevent a safe move, and the safe move is defined as O without justification (typically in the second player’s turn, playing O might not always be safe due to potential SOS threats from adjacent cells). - **Argument Structure:** The proof combines two potential strategies (traps and safe moves), but their integrations are vague. The safe move claim is not proven logically; it relies on an unverified implicit assumption, similar to the trapping strategy. This creates gaps in reasoning, and special-case analysis for safe moves cannot replace rigorous proof. - **Logical Fallacies:** The proof claims that traps ”kill” the opponent, but this is not demonstrated; it is asserted without showing that all first player responses are covered. The safe move relies on misstated bordering; for instance, corner cells have fewer neighbors, but the grid is linear, making the border condition ill-defined for edges. **(2) Computational Accuracy** - **Mathematical Calculations:** There are no significant calculations, so computational accuracy is not an issue, but this implies mathematical insight is missing. **(3) Structural Completeness** - **Full Coverage:** The proof does not address cases where the first player might prevent trap construction or exploit weaknesses in the safe move strategy early on. It assumes the second player can always build traps or perform the safe move, but no upper bound is established. The problem requires a winning strategy, but no complete game tree or position is analyzed. - **Proof Flow:** The proof jumps between non-verified claims and undefined terms. The safe move is presented as evidence for the main claim, but it is only part of the argument; the entire strategy is incomplete. **(4) Notational Rigor** - **Mathematical Notation:** The proof uses non-standard language like ”trap” and ambiguous descriptions such as ”bordered by two empty squares” applied to squares with an unclear number of neighbors. Terms like $\square$” and variable ”S” is not formally defined. - **Justification and Definitions:** Key concepts like SOS completion are not fully substantiated. The claim of a safe O move lacks proof of time-independent security (e.g., it might lead to SOS if the grid has other S around). The proof is logically flawed and incomplete, with gaps in reasoning and undefined elements. ### False
@@
We can observe from this example above that the original proof lacks necessary clarification to make the winning strategy viable and logically sound, relying on instinctive arguments when providing evidence for why the strategy leads to a winning situation. Our model is able to correctly discern this erroneous logic flow, as seen in the green highlighted part.
@@
Trying to provide a statistical analysis of which type of error can be better recognized by our ProofRM, we sample 97 QA pairs and manually categorize them into three types: misusing unjustified assumptions, incomplete logic flow, and making up unsubstantiated conclusions.
@@
The category comes from our idea that a logic step can be roughly modeled as a syllogism. So we categorize the errors based on the three key elements of a syllogism: premises, logic flow, and conclusion. The first error type corresponds to issues with premises, the second corresponds to issues with logic flow, and the third corresponds to issues with conclusions. We find most errors are due to incomplete logic flow, which indicates that the model fails to chain reasoning steps rigorously. The results are shown in Table 9.
@@
Let’s analyze the results in the table. Notice that the performance of the base model is even lower than a random guess (50%), suggesting the model easily believes incorrect proofs. ProofRM improves accuracy across these error types, with higher gains in the second and third categories, where the base models show weaker performance. It shows that our RL stage indeed helps our model to avoid the easy-believing behavior and find these errors.
@@ A.9 Ablation for balanced token weighting
In this section, we demonstrate the ablation results for balanced token weighting, giving a quantitative analysis on the resulting stability advantages caused by the introduction of this trick. We display the output length and overall reward of different token weighting strategies, namely GRPO-style, DAPO-style and balanced token weighting, throughout the training procedure in Figure 5 and Figure 6 respectively. The figures demonstrate the effectiveness of balanced token weighting. Although the reward for GRPO-style token weighting is increasing with training steps, upon closer inspection of the output we can see that the outputs contain overly short or even no reasoning, indicating the existence of behavior degeneration.
Figure 5: Output length comparison of different token weighting strategies: GRPO-style, DAPO-style and balanced token weighting.
Figure 6: Reward comparison of different token weighting strategies: GRPO-style, DAPO-style and balanced token weighting.
@@ A.10 Experiment results on other datasets
In this section we demonstrate the performance of our ProofRM on other evaluation sets, including both those readily provided by the community and those that we constructed ourselves. Table 10 displays the performance of ProofRM with differing model sizes on ProofBench (Ma et al., 2025). We can observe that our ProofRM outperforms baseline models of all sizes.
@@
Table 11: Results of ProofRM on mathematical proofs of different styles (forum, textbook, research article).
@@ A.11 Ablation on data diversity dimensions
Table 12: Ablation results on different data diversity dimensions. w/o Multi-sourced Dataset indicates that only samples from OlympiadBench are kept. w/o Extended LLM Usage indicates that only samples from Deepseek-R1 are kept, w/o Diverse Prompt Method indicates that only samples using rephrase are kept. Here the accuracy is evaluated on the our test set.
@@
We may observe from Table 12 that each dimension contributes to the diverse data quantity and also the final performance. Specifically, the diversity introduced by the ”prompt method” is the key to train a generalized model, where the model without this diversity shows worse accuracy than its base (69.9% v.s. 71.6%). The results also indicate the necessity and advantages of scaling up the dataset, further justifying our pipeline to produce data in a scalable way.